# FedBChain: A Blockchain-enabled Federated Learning Framework for Improving DeepConvLSTM with Comparative Strategy Insights*

### v1

@@ I INTRODUCTION
However, this was doubted and reassessed by [7], whose study used a single LSTM layer in a recurrent layer structure for model training. The results showed that the DeepConvLSTM structure based on a two-layer LSTM was outperformed by the single-layer DeepConvLSTM (with the LSTM module of the second layer removed) in terms of accuracy, recall, and F1-scores for four of the five datasets. Specifically, [7] achieved an average of 62% fewer trainable parameters across the architecture with this adaptation and reduced the overall training time by 48%. Although [7] achieved remarkable success on the DeepConvLSTM architecture based on single-layer LSTM, it is only applicable to monolithic servers, and in the large-scale distributed training environments, even if the performance of monolithic architectures is optimized again, it cannot effectively stifle the security and privacy issues of personal data. Therefore, in order to better handle the privacy and security of user data during training, federated learning [8] is proposed as an effective solution in large-scale distributed training environments. Federated learning is a decentralized machine learning technique that allows collaborative model training among multiple data owners while the data is always stored locally at the data owner.
Therefore, this project develops FedBChain, a federated learning framework specialized for large-scale distributed training environments, on top of DeepConvLSTM. This framework employs distributed training instead of the centralized training approach, thus enhancing data privacy and security. In addition, compared to the general federated learning process, FedBChain employs blockchain technology [9] as the global model in designing the global model. Generally, federated learning usually relies on a central node for aggregation and distribution of model parameters, which may have a single point of failure that jeopardizes the overall system performance, therefore, the design of this framework uses blockchain technology instead of a centralized global model. Meanwhile, the deployment and execution of smart contracts automatically performs model aggregation and distribution of the latest parameters to further ensure the robustness and availability of the system.
@@
  • This project introduces the FedBChain framework, which is based on a single-layer DeepConvLSTM architecture for federated learning training approach to ensure that data security and privacy are enhanced throughout the training process. The framework is suitable for large-scale distributed federated training.

### v2

@@ I INTRODUCTION
However, this was doubted and reassessed by [7], whose study used a single LSTM layer in a recurrent layer structure for model training. The results showed that the DeepConvLSTM structure based on a two-layer LSTM was outperformed by the single-layer DeepConvLSTM (with the LSTM module of the second layer removed) in terms of accuracy, recall, and F1-scores for four of the five datasets. Specifically, [7] achieved an average of 62% using fewer trainable parameters across the architecture with this adaptation and reduced the overall training time by 48%. Although [7] achieved remarkable success on the DeepConvLSTM architecture based on single-layer LSTM, it is only applicable to monolithic servers, and in the large-scale distributed training environments, even if the performance of monolithic architectures is optimized again, it can not effectively stifle the security and privacy issues of personal data. Therefore, in order to better handle the privacy and security of user data during training process, federated learning [8] is proposed as an effective solution in large-scale distributed training environments. Federated learning is a decentralized machine learning approach that enables collaborative model training across multiple data owners, ensuring that data remains stored locally with each owner.
Therefore, this project develops FedBChain, a federated learning framework specialized for large-scale distributed training environments, on top of DeepConvLSTM. This framework employs distributed training approach instead of the centralized training approach, thus enhancing data privacy and security. In addition, compared to the general federated learning process, FedBChain employs blockchain technology [9] as the global model in designing the global model. Generally, federated learning usually relies on a central node for aggregation and distribution of model parameters, which may have a single point of failure that risks the overall system performance, therefore, the design of this framework uses blockchain technology instead of a centralized global model. Meanwhile, the deployment and execution of smart contracts automatically performs model aggregation and distribution of the latest parameters to further ensure the robustness and availability of the system.
@@
  • This project introduces the FedBChain framework, which utilizes a single-layer DeepConvLSTM architecture for federated learning. This approach enhances data security and privacy throughout the training process. The framework is well-suited for large-scale distributed federated training.